OpenStudy (anonymous):

For the following rational function, list and label any x and y intercepts, as well as the vertical, horizontal, and oblique/slant asymptotes and holes (if any). Write the intercepts as coordinates; write asymptotes as equations. g(x)= [x^3 -4x]/[(x+2)(9x-1)]

OpenStudy (anonymous):

@badmood @CaptainFluffy

OpenStudy (anonymous):

@Hero @jigglypuff314 @Ashleyisakitty

OpenStudy (anonymous):

For starters, you can simplify your function, g: \[g(x)=\dfrac{x^3-4x}{(x+2)(9x-1)}=\dfrac{x(x-2)}{9x-1}\] Solving your question is a lot easier now.